# Multi-modal Retrieval Augmented Multi-modal Generation: Datasets, Evaluation Metrics and Strong Baselines

### v3

@@ Abstract
We present a systematic investigation of Multi-modal Retrieval Augmented Multi-modal Generation (M2RAG), a novel task that enables foundation models to process multi-modal web content and generate multi-modal responses, which exhibits better information density and readability. Despite its potential impact, M2RAG remains understudied, lacking comprehensive analysis and high-quality data resources. To address this gap, we establish a comprehensive benchmark through a rigorous data curation pipeline, and employ text-modal metrics and multi-modal metrics based on foundation models for evaluation. We further propose several strategies for foundation models to process M2RAG effectively and construct a training set by filtering high-quality samples using designed metrics. Our extensive experiments demonstrate the reliability of our proposed metrics, a landscape of model performance within our designed strategies, and show that our fine-tuned 7B-8B models outperform the state-of-the-art GPT-4o model. Additionally, we perform fine-grained analyses across diverse domains and validate the effectiveness of our designs in data curation pipeline. All resources, including codes, datasets, and model weights, will be publicly released.
[…]
tcb@breakable
Multi-modal Retrieval Augmented Multi-modal Generation:
Datasets, Evaluation Metrics and Strong Baselines
Zi-Ao Ma1, Tian Lan1, Rong-Cheng Tu2, Yong Hu3 Yu-Shi Zhu1, Tong Zhang1, Heyan Huang1, Xian-Ling Mao1thanks:  Corresponding author 1School of Computer Science and Technology, Beijing Institute of Technology, China 2Nanyang Technological University, Singapore, 3WeChat AI, Tencent Inc., China {maziaoylwt,lantiangmftby}@gmail.com, rongcheng.tu@ntu.edu.sg rightyonghu@tencent.com, maoxl@bit.edu.cn https://github.com/maziao/M2RAG
@@
1 Introduction
Retrieval Augmented Generation (RAG) Lan et al. (2023); Sun et al. (2024); Li et al. (2022) and its multi-modal extensions  Chen et al. (2022); Joshi et al. (2024); Ye et al. (2024) enhance foundation models by incorporating external knowledge and multi-modal data. While these methods improve response quality, they remain limited to textual outputs, which may fail to provide sufficient clarity in some scenarios, such as instructional guides or spatial reasoning. For instance, as shown in Figure 1, a text-only response explaining paper airplane folding steps may be difficult for users to follow without accompanying visual illustrations.
@@ 1 Introduction
As the saying goes, A single image is worth a thousand words. Visual elements can significantly enhance comprehension, especially in instructional and knowledge-intensive domains where textual descriptions alone may be insufficient. Motivated by this, the Multi-modal Retrieval-Augmented Multi-modal Generation (M2RAG) has been introduced  Zhu et al. (2024); Wu et al. (2023); Li et al. (2024), a novel task that retrieves and integrates multi-modal content to produce responses with a mixed layout of text and image. By integrating relevant images into textual responses, M2RAG enhances comprehension and usability across various application scenarios. As illustrated in Figure 1, M2RAG not only generates textual explanations for step-by-step folding a paper airplane but also seamlessly incorporates illustrative images where necessary, significantly improving user understanding. Compared to existing RAG methods, M2RAG introduces new challenges in multi-modal reasoning and synthesis. It requires foundation models to achieve a deeper understanding of multi-modal data, accurately capture content and cross-modal relationships, and generate coherent responses that seamlessly integrate text and images.
@@
However, given the early research stage, there is a lack of systematic analysis and high-quality data resources for the M2RAG task. To fill this gap, we make following contributions: (1) Benchmark: A comprehensive benchmark with 10 topics constructed by our proposed rigorous data curation pipeline; (2) Evaluation Metric: A suite of reliable text-only and multi-modal evaluation metrics based on Large Language Models (LLMs) and Multi-modal LLMs (MLLMs); (3) Generation Strategy: Two generation strategies for effectively tackling M2RAG task: single-stage and multi-stage approaches; and (4) Training Dataset: A training dataset constructed by filtering high-quality samples using our designed multi-modal evaluation metrics, which is used to improve the performance of 7B-8B LLMs and MLLMs.
@@
Based on our extensive experiments and automatic evaluation of advanced LLMs and MLLMs, we present four key findings: (1) Metric Reliability: The used evaluation metrics exhibit strong correlations with human judgments, even outperform inner-correlation of human annotators in evaluating image helpfulness dimension. This validates the reliability of our automatic evaluation framework; (2) Model Performance Analysis: Experimental results on 12 advanced LLMs and MLLMs provides valuable insights for M2RAG task. For example, LLMs consistently outperforms MLLMs, revealing MLLM’s limited multi-modal understanding and generation abilities. Besides, multi-stage strategy integrates more relevant images into responses, exhibiting better overall quality; (3) Effectiveness of Training Dataset: Fine-tuning 7B-8B LLMs and MLLMs on our curated dataset yields performance that exceeds GPT-4o, underscoring the quality and effectiveness of our training dataset; and (4) Data Curation Benefits: Ablation studies quantify cross-domain performance variations and verify the positive contributions of our designs in benchmark data curation pipeline, including detailed context for images and the inclusion of the auxiliary images. These observations and phenomena promote an in-depth understanding of M2RAG. We hope our data resources and these discoveries could spur future research in this field.
@@ 2 Related Work
RAG Li et al. (2022) has been widely used to improve the generation quality of language models by incorporating external knowledge Lan et al. (2023), which is usually retrieved by using BM25, dense retrieval model or search engine Lee et al. (2021); Asai et al. (2023). Recently, given the increasing capabilities of LLMs, there is an emerging way to simply concatenate all retrieved data into the context of LLMs for generation Sun et al. (2024).
@@
Despite the great potential of RAG, it can only handle textual inputs and cannot utilize the rich information in multi-modal data such as images and videos Su et al. (2023). While earlier efforts to incorporate multi-modal inputs for text generation often relied on carefully designed frameworks Xiao et al. (2024), a new paradigm has emerged that emphasizes pre-training MLLMs to accomplish this task more directly and efficiently Chen et al. (2024); Wang et al. (2024a); Yao et al. (2024); Shen et al. (2023); Riedler and Langer (2024), with representative models including GPT-4o Hurst et al. (2024) and Llama-3.2-Vision Dubey et al. (2024). While these multi-modal models greatly expand the range of LLMs’ applications, they remain limited to textual outputs, restricting their ability to deliver rich multi-modal responses to users Zhu et al. (2024).
@@
In real-world scenarios, humans naturally interact with multi-modal data, such as browsing web pages that combine text, images, and videos in mixed layouts Zhu et al. (2024). Consequently, it is crucial to develop foundation models that not only generate plain text responses to user queries but also incorporate relevant multi-modal data to enhance readability and user engagement. This approach embodies the principle—A single image is worth a thousand words, emphasizing the value of visual elements in effective communication. To the best of our knowledge, MExBERT Singh et al. (2021) took the first step in this research direction, which retrieves one image given the user query and generated response. Beyond retrieving the images, recent works also focus on generating the corresponding images for model-generated responses, like Next-GPT Wu et al. (2023), TextBind Li et al. (2024), Janus Wu et al. (2024) and Emu3 Wang et al. (2024b). Unlike these works, our study focuses on M2RAG task, which aims to dynamically select the multi-modal content from multiple multi-modal inputs (text or images) to construct final multi-modal responses without generating any visual elements. To the best of our knowledge, MuRAR Zhu et al. (2024) is the most closely related works to our study. However, they falls short of modeling multi-modal content holistically. For example, MuRAR generates initial responses as plain text based solely on user queries and retrieved textual content, without incorporating the corresponding multi-modal data. These approaches differs significantly from the setup of our proposed M2RAG task, which emphasizes the understanding the content and relationship among multi-modal input data.
@@ 3 Task Formulation
While previous studies have demonstrated the strong capabilities of LLMs to answer user queries grounded in retrieved documents, their inference time grows significantly with longer input sequences Sun et al. (2024). This challenge is more severe in the M2RAG task, where models must handle extensive textual and visual data across multiple documents. To address this challenge, we introduce the In-Doc Retrieval, which selects the most relevant and useful elements from $K$ to reduce inference costs. As illustrated in Figure 2 (Step 3), a retrieval model $M_{R}$ assesses the relevance of each element $E_{i,j}$ in the knowledge base $K$ with respect to the user query $Q$: $M_{R}(Q,K)$. The top-$k$ most relevant elements are selected to form a refined and concise knowledge base, $K_{\text{In-Doc}}$. Importantly, when a visual element is selected, its associated textual context is also retrieved to ensure coherence and provide richer input information for the generation process. The implementation details of the retrieval model is provided in Section 4.1.3.
@@
Figure 2: The overall framework of our proposed dataset construction and M2RAG pipeline. Step 1-3 represent the benchmark data curation pipeline and Step 4 demonstrates our proposed generation strategies.
@@ 4.1.1 Query Collection
The queries are sourced from the ELI5 Fan et al. (2019) dataset and undergo a series of filtering and classification procedures, as outlined in Step 1 of Figure 2. Specifically, we first exclude queries that do not require rich references and then discard those that do not necessitate visual information for answers. The filtered queries are then classified into 11 categories to ensure dataset balance and support further investigation of the proposed methods’ topic sensitivity. [^1] See Appendix B for details.
@@ 4.1.2 Data Preparation
Since the qualities of collected web pages and images are not guaranteed, we further propose Text Processing and Image Processing to post-process these multi-modal data to improve the quality of our dataset.
@@
Raw web pages often include irrelevant or redundant text, such as advertisements and web links, which must be removed to produce concise and relevant content for generation. The text processing pipeline consists of three steps: (1) Embedded Image Handling: Extract image URLs from web pages for image collection and replace them with placeholders in the text; (2) Text Cleaning: Remove text snippets matching predefined patterns, such as web links, to ensure cleaner and more relevant content; (3) Text Segmentation: Employ the text splitter Chase (2022) to split the cleaned text into smaller pieces, enabling efficient selection and processing.
@@
We conduct following steps to process images to ensure they are of high quality and relevant to the user query: (1) Caching & Conversion: All images are downloaded using their URLs and converted into widely accepted formats such as JPG, PNG, GIF, or WEBP. Images that cannot be successfully downloaded or converted are discarded; (2) Filtering: images smaller than a certain threshold or with a low CLIP-based Radford et al. (2021) similarity score to the query text are removed. Such images often consist of non-representative visual contents, such as icons, banners, etc. (3) Deduplication: duplicate or highly similar images are removed using PHash Zauner (2010) algorithm.
@@ 4.1.3 Element Evaluation
As described in Section 3, we need to perform In-Doc retrieval to select top-$k$ relevant text and image elements for efficient generation. To achieve this goal, we utilize DeepSeek-Chat DeepSeek-AI (2024) and MiniCPM-V-2.6 Yao et al. (2024) for evaluation, as shown in Figure 2 (Step 3). Our preliminary manual annotations reveal that these methods demonstrate strong correlations with human evaluations in measuring correlation and significantly outperforms the traditional embedding-based models like CLIP Radford et al. (2021), thereby resulting in better quality of our benchmark [^4] Please refer to Appendix C for more details..
In summary, given these data resources, we sample 75 queries for each category (other than Others), resulting in 750 queries for the benchmark dataset.
@@ 4.2 Evaluation Metrics
To examine the performance of LLMs and MLLMs on our proposed benchmark, we introduce four text-modal and four multi-modal metrics. Since the generated responses are open-ended, most evaluation metrics are implemented by prompting advanced GPT-4o model in a reference-free manner Lan et al. (2024b).
@@
Following prior works Zhu et al. (2024), we evaluate the quality of textual generations using both linguistic and RAG metrics Es et al. (2023); Lan et al. (2024a): (1) Fluency assesses the linguistic quality of model-generated text, ensuring the grammatical correctness, coherence, and readability Su et al. (2022); Es et al. (2023); (2) Relevance evaluates how well the model-generated textual content aligns with the given user query Lan et al. (2022); Es et al. (2023); (3) Context Precision measures the proportion of relevant chunks in $K_{\text{In-Doc}}$ by analyzing the overlap of their key concepts Es et al. (2023). (4) Faithfulness measures the accuracy of generations in representing information from $K_{\text{In-Doc}}$, focusing on factual alignment to avoid fabricated or misleading details Es et al. (2023); Sun et al. (2024).
@@
Unlike evaluation metrics that focus solely on image quality Tu et al. (2024), our work evaluates the interplay between text and images, covering four key aspects: (1) Image Coherence examines the logical and coherent alignment of images with their surrounding text, ensuring that the visual content enhances and complements the textual narrative Zhu et al. (2024); (2) Image Helpfulness evaluates the contributions of images to the user’s understanding of the text, assessing whether visuals provide additional insights, clarify complex ideas, or support textual details Zhu et al. (2024); (3) Image Reference verifies the appropriateness and alignment between images and their textual references. Inadequate or incorrect references of images lead to the confusion during reading; (4) Image Recall measures the proportion of highly relevant, informative, and important images incorporated into the generations. These crucial images are evaluated and collected by Element Evaluation (Section 4.1.3).
@@
We compute the overall score by taking the average of all text-modal and multi-modal metrics, reflecting the overall capabilities of models. Scores are scaled to [0, 100], with higher values indicating superior performance. For additional details on the specific prompts used for each metric, please refer to Appendix E.
@@ 4.3 Generation Strategy for M2RAG
Previous work MuRAR Zhu et al. (2024) generates and integrates text and images independently. We refer to this approach as the separate strategy, which neglects interactions between multi-modal data during text generation. To address this limitation, we propose two joint modeling strategies that explicitly capture the relationships between text and images: the single-stage strategy and the multi-stage strategy, as illustrated in Step 4 of Figure 2.
@@
where $G,Q,K_{\text{In-Doc}}$ denotes the task guidelines, user query and knowledge base. In the prompt $P$, each image is represented using Markdown format: ![IMAGE_CONTENT](PSEUDO_URL): (1) For MLLMs, IMAGE_CONTENT refers to encoded embeddings of the image, with the entire placeholder positioned where the image originally appears in the context Dubey et al. (2024); Wang et al. (2024a), ensuring coherence between the image and the surrounding text; (2) For LLMs, we convert the image into a detailed textual description as IMAGE_CONTENT, enabling the LLMs to comprehend the semantic information of the image. PSEUDO_URL serves as the identifier for each input image. In the output of the single-stage approach, images are also represented using the aforementioned Markdown format, where PSEUDO_URL indicates the index of the selected image.

### v4

@@ Abstract
We present a systematic investigation of Multi-modal Retrieval Augmented Multi-modal Generation (M2RAG), a novel task that enables foundation models to process multi-modal web content and generate multi-modal responses, which exhibits better information density and readability. Despite its potential impact, M2RAG remains understudied, lacking comprehensive analysis and high-quality data resources. To address this gap, we establish a comprehensive benchmark through a rigorous data curation pipeline, and employ text-modal metrics and multi-modal metrics based on foundation models for evaluation. We further propose several strategies for foundation models to process M2RAG task effectively and construct a training set by filtering high-quality samples using our designed metrics. Our extensive experiments demonstrate the reliability of our proposed metrics, a landscape of model performance within our designed strategies, and show that our fine-tuned 7B-8B models outperform the GPT-4o model and approach the state-of-the-art OpenAI o3-mini. Additionally, we perform fine-grained analyses across diverse domains and validate the effectiveness of our designs in data curation pipeline. All resources, including codes, datasets, and model weights, will be publicly released.
1 Introduction
Retrieval Augmented Generation (RAG) [15, 30, 19] and its multi-modal extensions  [2, 12, 38] enhance foundation models by incorporating external knowledge and multi-modal data. While these methods improve response quality, they remain limited to textual outputs, which may fail to provide sufficient clarity in some scenarios, such as instructional guides or spatial reasoning. For instance, as shown in Figure 1, a text-only response explaining paper airplane folding steps may be difficult for users to follow without accompanying visual illustrations.
@@ 1 Introduction
As the saying goes, A single image is worth a thousand words. Visual elements can significantly enhance comprehension, especially in instructional and knowledge-intensive domains where textual descriptions alone may be insufficient. Motivated by this, the Multi-modal Retrieval-Augmented Multi-modal Generation (M2RAG) has been introduced  [41, 35, 20], a novel task that retrieves and integrates multi-modal content to produce responses with a mixed layout of text and image. By integrating relevant images into textual responses, M2RAG enhances comprehension and usability across various application scenarios. As illustrated in Figure 1, M2RAG not only generates textual explanations for step-by-step folding a paper airplane but also seamlessly incorporates illustrative images where necessary, significantly improving user understanding. Compared to existing RAG methods, M2RAG introduces new challenges in multi-modal reasoning and synthesis. It requires models to achieve a deeper understanding of multi-modal data, accurately capture content and cross-modal relationships, and generate coherent responses that seamlessly integrate text and images [41].
@@
However, given the early research stage, there is a lack of systematic analysis and high-quality data for M2RAG task. To fill this gap, we make the following contributions: (1) Benchmark: A comprehensive benchmark with 10 topics constructed by our rigorous data curation pipeline; (2) Evaluation Metric: A suite of reliable text-only and multi-modal evaluation metrics based on Large Language Models (LLMs) and Multi-modal LLMs (MLLMs); (3) Generation Strategy: Two generation strategies for effectively tackling M2RAG task: single-stage and multi-stage approaches; and (4) Training Dataset: A training dataset constructed by filtering high-quality samples using our designed multi-modal evaluation metrics, which is used to improve the performance of 7B-8B LLMs and MLLMs.
@@
Based on our extensive experiments and automatic evaluation of advanced LLMs and MLLMs, we present four key findings: (1) Metric Reliability: The designed evaluation metrics exhibit strong correlations with human judgments, even outperform inner-correlation of human annotators in evaluating image helpfulness dimension. This validates the reliability of our automatic evaluation framework; (2) Model Performance Analysis: Experimental results on 12 advanced LLMs and MLLMs provides valuable insights for M2RAG task. For example, LLMs consistently outperforms MLLMs, revealing MLLM’s limited multi-modal understanding and generation abilities. Besides, multi-stage strategy integrates more relevant images into responses, exhibiting better overall quality; (3) Effectiveness of Training Dataset: Fine-tuning 7B-8B LLMs and MLLMs on our curated dataset yields performance that exceeds GPT-4o, underscoring the quality and effectiveness of our training dataset; and (4) Data Curation Benefits: Ablation studies quantify cross-domain performance variations and verify the positive contributions of our designs in benchmark data curation pipeline, including detailed context for images and the inclusion of the auxiliary images. These observations and phenomena promote an in-depth understanding of M2RAG. We hope our data resources and these discoveries could spur future research in this field.
@@ 2 Related Work
RAG [19] has been widely used to improve the generation quality of language models by incorporating external knowledge [15], which is usually retrieved by using BM25, dense retrieval model or search engine [18, 1]. Recently, given the increasing capabilities of LLMs, there is an emerging way to simply concatenate all retrieved data into the context of LLMs for generation [30].
@@
Despite the great potential of RAG, it can only handle textual inputs and cannot utilize the rich information in multi-modal data such as images and videos [29]. While earlier efforts to incorporate multi-modal inputs for text generation often relied on carefully designed frameworks [36], a new paradigm has emerged that emphasizes pre-training MLLMs to accomplish this task more directly and efficiently [3, 32, 37, 26, 25], with representative models including GPT-4o [11] and Llama-3.2-Vision [7]. While these multi-modal models greatly expand the range of LLMs’ applications, they remain limited to textual outputs, restricting their ability to deliver rich multi-modal responses to users [41].
@@
In real-world scenarios, humans naturally interact with multi-modal data, such as browsing web pages that combine text, images, and videos in mixed layouts [41]. Consequently, it is crucial to develop foundation models that not only generate plain text responses to user queries but also incorporate relevant multi-modal data to enhance readability and user engagement. This approach embodies the principle—A single image is worth a thousand words, emphasizing the value of visual elements in effective communication. To the best of our knowledge, MExBERT [27] took the first step in this research direction, which retrieves one image given the user query and generated response. Beyond retrieving the images, recent works also focus on generating the corresponding images for model-generated responses, like Next-GPT [35], TextBind [20], Janus [34] and Emu3 [33]. Unlike these works, our study focuses on M2RAG task, which aims to dynamically select the multi-modal content from multiple multi-modal inputs (text or images) to construct final multi-modal responses without generating any visual elements. To the best of our knowledge, MuRAR [41] is the most closely related works to our study. However, they falls short of modeling multi-modal content holistically. For example, MuRAR generates initial responses as plain text based solely on user queries and retrieved textual content, without incorporating the corresponding multi-modal data. These approaches differs significantly from the setup of our proposed M2RAG task, which emphasizes the understanding the content and relationship among multi-modal input data.
@@ 3 Task Formulation
While previous studies have demonstrated the strong capabilities of LLMs to answer user queries grounded in retrieved documents, their inference time grows significantly with longer input sequences [30]. This challenge is more severe in the M2RAG task, where models must handle extensive textual and visual data across multiple documents. To address this challenge, we introduce the In-Doc Retrieval, which selects the most relevant and useful elements from $K$ to reduce inference costs. As illustrated in Figure 2 (Step 3), a retrieval model $M_{R}$ assesses the relevance of each element $E_{i,j}$ in the knowledge base $K$ with respect to the user query $Q$: $M_{R}(Q,K)$. The top-$k$ most relevant elements are selected to form a refined and concise knowledge base, $K_{\text{In-Doc}}$. Importantly, when a visual element is selected, its associated textual context is also retrieved to ensure coherence and provide richer input information for the generation process. The implementation details of the retrieval model is provided in Section 4.1.3.
@@
Figure 2: The framework of our proposed dataset construction and M2RAG pipeline. Step 1-3 represent the data curation pipeline and Step 4 demonstrates our proposed generation strategies.
@@ 4.1.1 Query Collection
The queries are sourced from the ELI5 [6] dataset and undergo a series of filtering and classification procedures, as outlined in Step 1 of Figure 2. Specifically, we first exclude queries that do not require rich references and then discard those that do not necessitate visual information for answers. The filtered queries are then classified into 11 topics to ensure dataset balance and support further investigation of the proposed methods’ topic sensitivity. [^1] See Appendix C for details.
@@ 4.1.2 Data Preparation
Raw web pages often include irrelevant or redundant text, such as advertisements and web links, which must be removed to produce concise and relevant content for generation. The text processing pipeline consists of three steps: (1) Embedded Image Handling: Extract image URLs from web pages for image collection and replace them with placeholders in the text; (2) Text Cleaning: Remove text snippets matching predefined patterns, such as web links, to ensure cleaner and more relevant content; (3) Text Segmentation: Employ the text splitter to split the cleaned text into smaller pieces, enabling efficient selection and processing.
@@
We conduct following steps to process images to ensure they are of high quality and relevant to the user query: (1) Caching & Conversion: All images are downloaded using their URLs and converted into widely accepted formats such as JPG, PNG, GIF, or WEBP. Images that cannot be successfully downloaded or converted are discarded; (2) Filtering: images smaller than a certain threshold or with a low CLIP-based [23] similarity score to the query text are removed. Such images often consist of non-representative visual contents, such as icons, banners, etc. (3) Deduplication: duplicate or highly similar images are removed using PHash [39] algorithm.
@@ 4.1.3 Element Evaluation
As described in Section 3, we need to perform In-Doc retrieval to select top-$k$ relevant text and image elements for efficient generation. To achieve this goal, we utilize DeepSeek-Chat [21] and MiniCPM-V-2.6 [37] for evaluation, as shown in Figure 2 (Step 3). Our preliminary manual annotations reveal that these methods demonstrate strong correlations with human evaluations in measuring correlation and significantly outperforms the traditional embedding-based models like CLIP [23], thereby resulting in better quality of our benchmark [^4] Please refer to Appendix D for more details..
In summary, given these data resources, we sample 100 queries for each category (other than Others), resulting in 1,000 queries for the benchmark dataset.
@@ 4.2 Evaluation Metrics
To examine the performance of LLMs and MLLMs on our proposed benchmark, we introduce four text-modal and four multi-modal metrics. Since the generated responses are open-ended, most evaluation metrics are implemented by prompting advanced GPT-4o model in a reference-free manner [17].
@@
Following prior works [41], we evaluate the quality of textual generations using both linguistic and RAG metrics [5, 16]: (1) Fluency assesses the linguistic quality of model-generated text, ensuring the grammatical correctness, coherence, and readability [28, 5]; (2) Relevance evaluates how well the model-generated textual content aligns with the given user query [14, 5]; (3) Context Precision measures the proportion of relevant chunks in $K_{\text{In-Doc}}$ by analyzing the overlap of their key concepts [5]. (4) Faithfulness measures the accuracy of generations in representing information from $K_{\text{In-Doc}}$, focusing on factual alignment to avoid fabricated or misleading details [5, 30].
@@
Unlike evaluation metrics that focus solely on image quality [31], our work evaluates the interplay between text and images, covering four key aspects: (1) Image Coherence examines the logical and coherent alignment of images with their surrounding text, ensuring that the visual content enhances and complements the textual narrative [41]; (2) Image Helpfulness evaluates the contributions of images to the user’s understanding of the text, assessing whether visuals provide additional insights, clarify complex ideas, or support textual details [41]; (3) Image Reference verifies the appropriateness and alignment between images and their textual references. Inadequate or incorrect references of images lead to the confusion during reading; (4) Image Recall measures the proportion of highly relevant, informative, and important images incorporated into the generations. These crucial images are evaluated and collected by Element Evaluation (Section 4.1.3).
@@
We compute the overall score by taking the average of all text-modal and multi-modal metrics, reflecting the overall capabilities of models. Scores are scaled to [0, 100], with higher values indicating superior performance. For additional details on the specific prompts used for each metric, please refer to Appendix F.
@@ 4.3 Generation Strategy for M2RAG
Previous work MuRAR [41] generates and integrates text and images independently. We refer to this approach as the separate strategy, which neglects interactions between multi-modal data during text generation. To address this limitation, we propose two joint modeling strategies that explicitly capture the relationships between text and images: the single-stage strategy and the multi-stage strategy, as illustrated in Step 4 of Figure 2.
@@
where $G,Q,K_{\text{In-Doc}}$ denotes the task guidelines, user query and knowledge base. In the prompt $P$, each image is represented using Markdown format: ![IMAGE_CONTENT](PSEUDO_URL): (1) For MLLMs, IMAGE_CONTENT refers to encoded embeddings of the image, with the entire placeholder positioned where the image originally appears in the context [7, 32], ensuring coherence between the image and the surrounding text; (2) For LLMs, we convert the image into a detailed textual description as IMAGE_CONTENT, enabling the LLMs to comprehend the semantic information of the image. PSEUDO_URL serves as the identifier for each input image. In the output of the single-stage approach, images are also represented using the aforementioned Markdown format, where PSEUDO_URL indicates the index of the selected image.